How much 12% alcohol solution must be mixed with 30% alcohol solution to get 60 liters of 27% alcohol solution?

Answer by checkley79(3341)

.30x+.12(60-x)=.27*60
.30x+7.2-.12x=16.2
.18x=16.2-7.2
.18x=9
x=9/.18
x=50 liters of 30% alcohol is used.
60-50=10 liters of 12% is used.
Proof:
.30*50+.12*10=.27*60
15+1.2=16.2
16.2=16.2